A 10.0 L gas cylinder has been filled with 6.20 moles of gas. You measure the pressure to be 3.50 atm . What is the temperature inside the tank?
Answer #1

Answer o gas Cylinder Volume = v= 10 Lit moles=n=6- 20 moles pressure=p=3.50 atm constant R = 0.0821 Litats mot K. TemperaturTe coa = 68-69K Temperature = 68.89k